OpenStudy (anonymous):

Find dy/dx of xy=cot^3(y+x)^2

OpenStudy (anonymous):

Implicit D!

OpenStudy (anonymous):

You need: what is the derivative of \[\frac{d}{dx}\cot(x)\]

OpenStudy (anonymous):

-csc^2(x)

OpenStudy (anonymous):

Ok so what is \[\frac{d}{dx}xy\] using implicit D? You need the product rule here.

OpenStudy (anonymous):

x+y?

OpenStudy (anonymous):

well the product rule gives you (ab)' = ab' + a'b

OpenStudy (anonymous):

You can use the chain rule on (x+y)^2 too.

OpenStudy (anonymous):

x+y

OpenStudy (anonymous):

Oh some angle identity, eh?

OpenStudy (anonymous):

ok, so implicit D would have \[\frac{d}{dx}xy = x\frac{d}{dx}y + y\frac{d}{dx}x\] which would then be just \[\frac{d}{dx}xy = x\frac{d}{dx}y + y\] or in short form \[\frac{d}{dx}xy = xy^{\prime} + y\]

OpenStudy (anonymous):

Implicit D is actually the chain rule but with variables. We don't know what y(x).